### v1

@@ 10 Limitations and Future Research
HeadZoom advances the field of spatial interaction by enabling fluid, continuous panning and zooming through natural head movement. Although promising, this approach also introduces new challenges and opportunities. HeadZoom demonstrates the feasibility of hands-free head-based interaction for 2D content navigation, but several limitations remain. First, the system relies solely on head pose as input, which constrains the granularity of control and may not scale well to tasks that require high-precision manipulation or sustained focus over extended sessions. In the future it would be good to explore other input alternatives, especially multimodal options, such as combining eye-gaze and head poose.
Another limitation is that the currently implemented system assumes stable seating posture and uniform viewing distance, which may not hold in ambulatory or multi-user VR contexts. In the future we should conduct experiments supporting a wider range of physical movement.
Finally, the research was conducted with relatively simple tasks. In order for the results to generalize, future studies should be conducted with a wider range of tasks. This will help establish how applicable technique is in a wider range of VR settings.
Future Research should address the following directions:
Ergonomics and Fatigue: Prolonged head movement may cause neck strain or discomfort for the user. Investigating ergonomic thresholds and incorporating adaptive mechanisms could mitigate fatigue and improve long-term usability [40].
Precision and Responsiveness: Head-based input must support fine-grained control and minimal latency for effective navigation. Further studies should examine how system responsiveness affects task accuracy and user satisfaction [15].
Multimodal Integration: Combining head-based control with complementary modalities, e.g. eye gaze or speech, may improve precision, reduce workload, and support diverse user needs [49].
Application-Specific Optimization: Future implementations should tailor head-based navigation to domain-specific requirements, including image granularity in medical imaging, map scale in GIS, and browsing patterns in photography [8].
By addressing these directions, future systems can harness the full potential of head-based interaction while ensuring usability, comfort, and domain adaptability.
@@ 11 Conclusions
This paper presented HeadZoom, a hands-free interaction technique that enables users to pan and zoom two-dimensional imagery using natural head motion. Through a controlled user study involving three interaction conditions: static, parallel, and tilt, we demonstrated that head-based control can be effective and usable in diverse image navigation tasks.
Among the tested methods, the Parallel mode consistently minimized physical effort while maintaining task performance and was rated most favorably by participants. Statistical analysis confirmed significant reductions in head movement and zoom-related actions, suggesting its suitability for applications requiring sustained or precise visual exploration. Although the Static condition yielded slightly faster average times, it imposed greater physical demand, particularly in rotational metrics. Although promising in theory, the Tilt condition was often perceived as disorienting due to its dynamic image realignment.
Our findings underscore the feasibility of replacing traditional input modalities with controller-free embodied techniques in mixed reality contexts. Future research should explore multi-modal extensions (e.g., gaze, speech) and domain-specific adaptations for clinical, cartographic, and creative workflows. In addition, more research is needed on adaptive calibration strategies and long-term ergonomics to enhance robustness and inclusivity among broader user populations.
Supplemental Materials
Supplemental materials include:
The questionnaires filled by Subjects during the experiments
a copy of all the images used during the user tests. The dataset includes seven images. One was used for training (resized_city) and the remaining six were used for training.
a companion video
These are available as additional items in the precisionconference submission
@@ Acknowledgements.
The authors wish to thank A, B, and C. This work was supported in part by a grant from XYZ.References
@@
- [13] A. Cockburn and A. Karlson. A review of zoomable user interfaces: Foundations and future directions. ACM Computing Surveys, 40(2):5:1–5:39, 2008. doi: 10.1145/1357054.1357213
@@
- [20] C. Harrison and A. Dey. Lean and zoom: Proximity-aware user interface for touchscreen devices. In Proceedings of CHI, pp. 1003–1012, 2008. doi: 10.1145/1357054.1357213

### v2

@@ 10 Limitations and Future Research
HeadZoom advances the field of spatial interaction by enabling fluid, continuous panning and zooming through natural head movement. Although promising, this approach also introduces new challenges and opportunities. HeadZoom demonstrates the feasibility of hands-free head-based interaction for navigating 2D content, but several limitations remain. First, the system relies solely on head pose as input, which constrains the granularity of control and may not scale well to tasks that require high-precision manipulation or sustained focus over extended sessions. In the future it would be good to explore other input alternatives, especially multimodal options, such as combining eye-gaze and head pose.
Another limitation is that the current system assumes stable seating posture and uniform viewing distance, which may not hold in ambulatory or multi-user VR contexts. In the future we should conduct experiments supporting a wider range of physical movement.
Finally, the research was conducted with relatively simple tasks. In order for the results to generalize, future studies should be performed with a broader variety of tasks. This will help establish the applicability of the technique in a wider range of VR settings.
Future Research should address the following limitations and issues: Ergonomics and Fatigue: Prolonged head movement may cause neck strain or discomfort for the user. Investigating ergonomic thresholds and incorporating adaptive mechanisms could mitigate fatigue and improve long-term usability [48]. Precision and Responsiveness: Head-based input must support fine-grained control and minimal latency for effective navigation. Further studies should examine how system responsiveness affects task accuracy and user satisfaction [19]. Multimodal Integration: Combining head-based control with complementary modalities, e.g. eye gaze or speech, may improve precision, reduce workload, and support diverse user needs [61]. Application-Specific Optimization: Future implementations should tailor head-based navigation to domain-specific requirements, including image granularity in medical imaging, map scale in GIS, and browsing patterns in photography [10].
@@ 11 Conclusions
This paper presented HeadZoom, a hands-free interaction technique that enables users to pan and zoom two-dimensional imagery using natural head motion. Through a controlled user study with three interaction conditions: static, parallel, and tilt, we found that head-based control can be effective in diverse image navigation tasks.
Among the tested methods, the Parallel mode minimized physical effort while maintaining task performance and was rated most favorably. Statistical analysis confirmed significant reductions in head movement and zoom-related actions, suggesting its suitability for applications requiring sustained or precise visual exploration. Although the Static condition yielded slightly faster average times, it imposed greater physical demand, particularly in rotational metrics. Although promising in theory, the Tilt condition was often perceived as disorienting due to its dynamic image realignment.
Our findings underscore the feasibility of replacing traditional input modalities with controller-free embodied techniques in MR. Future research should explore multi-modal extensions (e.g., gaze, speech) and domain-specific adaptations for clinical, cartographic, and creative workflows. In addition, more research is needed on adaptive calibration strategies and long-term ergonomics to enhance robustness and inclusivity among broader user populations.
@@ Acknowledgements.
The authors wish to thank Anderson Maciel for his comments, the UNESCO Chair on AI & VR and support from Fundação para a Ciência e a Tecnologia grants DOI:10.54499/UIDB/50021/2020, and 2022.09212.PTDC.References
- [1] H. Afkari, S. Eivazi, R. Bednarik, and S. Mäkelä. The potentials for hands-free interaction in micro-neurosurgery. In Proceedings of the 8th Nordic Conference on Human-Computer Interaction: Fun, Fast, Foundational, NordiCHI ’14, p. 401–410. ACM, 2014. doi: 10.1145/2639189.2639213
- [2] R. Atienza, R. Blonna, M. I. Saludares, J. Casimiro, and V. Fuentes. Interaction techniques using head gaze for virtual reality. In 2016 IEEE Region 10 Symposium (TENSYMP), pp. 110–114, 2016. doi: 10.1109/TENCONSpring.2016.7519387
@@
- [14] R.-C. Chang, C.-S. Hung, B.-Y. Chen, D. Jain, and A. Guo. Soundshift: Exploring sound manipulations for accessible mixed-reality awareness. In Proceedings of the ACM Designing Interactive Systems Conference, p. 116–132. ACM, 2024. doi: 10.1145/3643834.3661556
@@
- [16] A. Cockburn, A. Karlson, and B. B. Bederson. A review of overview+detail, zooming, and focus+context interfaces. ACM Comput. Surv., 41(1), Jan. 2009. doi: 10.1145/1456650.1456652
- [17] N. Correia, L. Alves, H. Correia, L. Romero, C. Morgado, L. Soares, J. C. Cunha, T. Romão, A. E. Dias, and J. A. Jorge. Instory. In Proceedings of the 2005 ACM ACE, 2005.
@@
- [24] C. Harrison and A. K. Dey. Lean and zoom: proximity-aware user interface and content magnification. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems, CHI ’08, p. 507–510. Association for Computing Machinery, New York, NY, USA, 2008. doi: 10.1145/1357054.1357135
@@
- [26] B. J. Hou, J. Newn, L. Sidenmark, A. Ahmad Khan, P. Bækgaard, and H. Gellersen. Classifying head movements to separate head-gaze and head gestures as distinct modes of input. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems, CHI ’23. ACM, 2023. doi: 10.1145/3544548.3581201
@@
- [31] H. Javaheri, O. Ghamarnejad, P. Lukowicz, G. A. Stavrou, and J. Karolus. Aras: Llm-supported augmented reality assistance system for pancreatic surgery. In Companion of the 2024 on ACM International Joint Conference on Pervasive and Ubiquitous Computing, UbiComp ’24, p. 176–180. ACM, 2024. doi: 10.1145/3675094.3677543
@@
- [42] P. Monteiro, H. Coelho, G. Gonçalves, M. Melo, and M. Bessa. Exploring the user experience of hands-free vr interaction methods during a fitts’ task. Computers & Graphics, 117:1–12, 2023. doi: 10.1016/j.cag.2023.10.005
- [43] P. Monteiro, G. Gonçalves, H. Coelho, M. Melo, and M. Bessa. Hands-free interaction in immersive virtual reality: A systematic review. IEEE Transactions on Visualization and Computer Graphics, 27(5):2702–2713, 2021. doi: 10.1109/TVCG.2021.3067687
@@
- [53] Y. Y. Qian and R. J. Teather. The eyes don’t have it: an empirical comparison of head-based and eye-based selection in virtual reality. In Proceedings of the 5th Symposium on Spatial User Interaction, SUI ’17, p. 91–98. ACM, 2017. doi: 10.1145/3131277.3132182
- [54] S. F. Rauh, C. Bogdan, G. Meixner, and A. Matviienko. Navigating the virtuality-reality clash: Reflection and design patterns for industrial mixed reality applications. In Proceedings of the 2024 ACM Designing Interactive Systems Conference, DIS ’24, p. 2247–2266. ACM, 2024. doi: 10.1145/3643834.3660700
@@
- [56] D. Rudi, I. Giannopoulos, P. Kiefer, C. Peier, and M. Raubal. Interacting with maps on optical head-mounted displays. In Proceedings of the 2016 Symposium on Spatial User Interaction, SUI ’16, p. 3–12. ACM, 2016. doi: 10.1145/2983310.2985747
@@
- [59] D. Schott, M. Kunz, F. Heinrich, J. Mandel, A. Albrecht, R. Braun-Dullaeus, and C. Hansen. Stand alone or stay together: An in-situ experiment of mixed-reality applications in embryonic anatomy education. In Proceedings of the 30th ACM Symposium on Virtual Reality Software and Technology, VRST ’24. ACM, 2024. doi: 10.1145/3641825.3687706
@@
- [66] B. V. Syiem, S. Webber, R. M. Kelly, Q. Zhou, J. Goncalves, and E. Velloso. Augmented reality at zoo exhibits: A design framework for enhancing the zoo experience. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems, CHI ’24. ACM, 2024. doi: 10.1145/3613904.3642015
- [67] M. Vojtechovska, S. Popelka, and P. Kubíček. Gaze controlled maps: scoping review of gaze-based interactions in geovisualisations. International Journal of Digital Earth, 18(1):2510563, 2025.
- [68] H. Wang, L. Sidenmark, F. Weidner, J. Newn, and H. Gellersen. Headshift: Head pointing with dynamic control-display gain. ACM Trans. Comput.-Hum. Interact., 32(1), Apr. 2025. doi: 10.1145/3689434
@@
- [71] Y. Yan, Y. Shi, C. Yu, and Y. Shi. Headcross: Exploring head-based crossing selection on head-mounted displays. Proc. of ACM on Interactactive, Mobile, Wearable and Ubiquitous Technologies, 4(1), Mar. 2020. doi: 10.1145/3380983
- [72] Y. Yan, C. Yu, X. Yi, and Y. Shi. Headgesture: Hands-free input approach leveraging head movements for hmd devices. Proc. of ACM on Interactactive, Mobile, Wearable and Ubiquitous Technologies, 2(4), Dec. 2018. doi: 10.1145/3287076